OpenStudy (anonymous):

If the sum of interior angle measures of a polygon is 900, how many sides does the polygon have?

OpenStudy (alexandervonhumboldt2):

use formula: 180*(n-2)=900. where n is number of sides.

OpenStudy (alexandervonhumboldt2):

n-2=900/180 n=7

OpenStudy (anonymous):

Thanks! Thats the answer I got I just wasnt forsure.